# Quantum Mechanics Relative to a Quantum Reference System: a Relative State Approach

### v2

@@ V Recovering the Schrödinger Equation
The evolution equations (27) or (28) plays a role analogous to that of the standard Schrödinger equation in describing the evolution of quantum states with respect to absolute external parameters. Therefore, in the following, we will start from the equation (27) and make some approximations to return to the Schrödinger equation limit. First, we linearize the Ricci curvature as $R_{\bar{i}j}\approx-\frac{1}{2}\Delta_{x}h_{\bar{i}j}$, where $\Delta_{x}$ is a linearized the Laplacian-Beltrami operator with respect to $x_{i}$, then eq.(27) becomes
@@
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}=\frac{\partial}{\partial t_{a}}\langle X_{i}|X_{k}\rangle=\frac{\partial\langle X_{i}|}{\partial t_{a}}|X_{k}\rangle+\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}\approx\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}=e^{-i\omega_{a}t^{a}}\langle X_{i}|\left(\frac{\partial}{\partial t^{a}}-i\omega_{a}\right)|e_{k}\rangle$
@@
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}\approx\left(-i\omega^{a}\right)h_{\bar{i}k}$
@@
To sum up, we have demonstrated that the equation (27) or its linearized approximations (31) or (32) describe the evolution of the metric $h_{\bar{i}j}$ or the basis vectors $|X_{i}\rangle$ in the subspace $\mathscr{H}_{X}$ with respect to the entangled state vector $|T\rangle$ in $\mathscr{H}_{T}$. This is because we observe that the slow-changing part $|e_{i}\rangle$ of the basis vectors $|X_{i}\rangle$ satisfies an approximate “free” Schrödinger evolution equation, where the fast-changing characteristic frequency $\omega^{a}$ a plays a role analogous to the mass $M=\hbar\omega^{a}$, in which the mass and the Planck’s constant emerge, and the potential term in the Schrödinger equation is embodied in the connection (scalar extrinsic curvature $K_{a}$) of the covariant derivative.
@@ VI Non-Inertial Effects
where $F_{\bar{a}b}$ reflects a kind of "inertial force" that arises due to the choice of quantum reference system when $|X\rangle$ is considered relative to $|T\rangle$. Under the approximation (34), we return to the situation of an approximately inertial frame where it is inertial-force-free.
@@
in which $G_{\bar{a}b}$ and $\Omega_{\bar{a}b}$ are the real and imaginary part of the metric. As a metric coming from the comparison between $|X^{i}\rangle$ and $|T^{a}\rangle$, it can be interconverted with the previous metrics $s_{\bar{a}b}$ and $h_{\bar{i}j}$ through a tensor coordinate transformation, for example,
$g_{\bar{a}b}=\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X^{j}}{\partial t^{b}}\right\rangle h_{i\bar{j}}=\left\langle\frac{\partial T_{a}}{\partial x_{i}}\right|\left.\frac{\partial T_{b}}{\partial x_{j}}\right\rangle h_{i\bar{j}}=\left\langle\frac{\partial T_{a}}{\partial t_{c}}\right|\left.\frac{\partial T_{b}}{\partial t_{d}}\right\rangle s_{\bar{c}d}$ (45)
@@
$\langle X^{i}(t+dt)|X_{i}(t)\rangle=\langle X^{i}(t)|X_{i}(t)\rangle+\frac{1}{2}\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle dt^{a*}dt^{b}+...=e^{d\varTheta}=1+d\varTheta+$... (47)
@@
$\textrm{Re}\left(d\varTheta\right)$ $=\frac{1}{2}\textrm{Re}\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle dt^{a*}dt^{b}=\frac{1}{2}G_{\bar{a}b}dt^{a*}dt^{b}$
$=\left\|\langle X^{i}(t)|X_{i}(t)\rangle-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|=\left\|1-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|\in[0,1]$ (48)
@@
Comparing this equation with the definition $\Omega_{\bar{a}b}\equiv g_{\bar{a}b}-g_{\bar{a}b}^{*}$ by using (46), we conclude that $\Omega_{\bar{a}b}$ can be interpreted as the Berry curvature (or field strength) $B_{\bar{a}b}$
@@
Next, let’s discuss some physical implications of the real and imaginary parts of the non-integrable phase $\varTheta$. We observe that $\textrm{Re}\left(d\varTheta\right)=\left\|1-\langle X^{i}(t+dt)|X_{i}(t)\rangle\right\|=\frac{1}{2}G_{\bar{a}b}dt^{a*}dt^{b}$ can also be interpreted as an infinitesimal stretching of the length of the state vector $|X\rangle$ (i.e. $\left\||X\rangle\right\|$), which makes the evolution of $|X\rangle$ with respect to $|T\rangle$ no longer unitary in general. Consequently, previously we could only derive the Schrödinger equation with unitary evolution under the assumption of slow-changing of the scalar extrinsic curvature (34) $G_{\bar{a}b}=\textrm{Re}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)\approx 0$. This is also an inevitable consequence of the relative evolution of subsystems in entangled states, resulting from the quantum superposition (quantum second-moment spreading (Luo2015Dark, ; Luo:2023eqf, ; 2024arXiv240809630L, )) of the $T$ subsystem that serves as the reference frame.
This breakdown of unitarity can also be interpreted as an effect of additional second-order fluctuations or extra broadening. Considering the second-moment spreading of the effective energy $E$ corresponding to the second-moment fluctuations of $t$, namely $\langle\delta E^{2}\rangle=\langle E^{2}\rangle-\langle E\rangle^{2}$, it leads to $\langle X^{i}(t+dt)|X_{i}(t)\rangle=1-\frac{1}{2}\langle\delta E^{2}\rangle dt^{2}$. Therefore, the metric $g_{\bar{a}b}$ is directly determined by the second-moment quantum fluctuations of $t$ or the second-moment fluctuations of the effective energy, i.e.
$g_{\bar{a}b}dt^{a*}dt^{b}\sim\langle\delta E^{2}\rangle|dt|^{2}$ (55)
This is nothing but the Anandan-Aharonov theorem (Anandan:1990fq, ). Hence, this “inertial force” $F_{\bar{a}b}=g_{\bar{a}b}\sim\langle\delta E^{2}\rangle$ can also be regarded as arising from the quantum fluctuations of energy or quantum fluctuations of the clock reference frame. We also observe that the breakdown of unitarity caused by the second-order fluctuations of the quantum clock and the broadening of spectral lines (Luo:2023eqf, ) both are coming from the additional “inertial force” effect. Conversely, the effect of the “inertial force” always seems to be associated with changes in the norm of the state vector or the breakdown of quantum unitarity. This non-unitarity resulting from the broadening of the quantum clock or spectral lines will be generalized to more general quantum spacetime reference frames in (Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ; 2024arXiv240809630L, ), and is the fundamental reason of coordinate transformation anomalies, the cosmological constant, and gravity in general quantum spacetime reference frames.
@@
If the phase $\varTheta=\frac{1}{2}\int_{c}dt^{a}K_{a}$ is integrable (independent to the path $c$), meaning that the local Berry curvature $B=B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ is closed ($dB=d^{2}K=0$), and so if this local Berry curvature corresponds to the imaginary part $\Omega_{\bar{a}b}$ of the local metric, the quantum state space equipped with this Hermitian complex metric must be a Kähler space. In this case, the gauge field strength 2-form $\frac{i}{2}B_{\bar{a}b}dt^{a*}\wedge dt^{b}\doteq\frac{i}{2}\Omega_{\bar{a}b}dt^{a*}\wedge dt^{b}$ serves as the Kähler form, reflecting the symplectic structure of the Kähler space. A Kähler space is one that simultaneously possesses a Riemannian structure, a symplectic structure, and is compatible with its complex structure. When the quantum state space is a Kähler space, or equivalently, the phase $\varTheta$ is integrable, due to the enhanced symmetry of the Kähler metric, which can be represented by a single Kähler potential, many problems by solving the multi-component Ricci-flat Kähler-Einstein equations can be reduced to solving a single complex Monge-Ampère equation, which simplifies the problem of solving the Kähler-Einstein equations. When there exists a constant state space volume form $\sqrt{\det h_{\bar{i}j}}$, i.e. a constant Kähler potential $\ln\sqrt{\det h_{\bar{i}j}}$, this degenerates to the case of a unitary quantum mechanics, so the state space of the unitary standard quantum mechanics is Kähler, which is suggested by literature e.g. (Ashtekar:1997ud, ; Bengtsson:2006rfv, ). In particular, when the curvature of its subspace is proportional to the metric, $R_{\bar{a}b}=(N+1)g_{\bar{a}b}$, the quantum state space is homeomorphic to the maximally symmetric sphere $S^{2N+1}$ and this Kähler metric $g_{\bar{a}b}$ is known as the Fubini-Study metric (Pandya:2006mrv, ; Matassa:2020dnc, ). However, the subspace metric $h_{\bar{i}j}$ of interest to us generally does not possess constant curvature or a constant local volume form, so the subspace metric of the quantum state here is generally not Fubini-Study in this framework. Nonetheless, in general, our phase $\varTheta$ is not strictly integrable (depending on the path and history), and the Berry curvature can no longer be locally defined. In this case, the quantum state subspace of the framework is not strictly Kähler but can be regarded as an almost Kähler manifold.
@@ VII Comparison with Some Past Works
Our paper achieves results that are similar in many respects to those in the work by Page and Wootters (PhysRevD.27.2885, ). However, there are also numerous differences between our paper and their study. Firstly, in Page and Wootters’ article, the Hamiltonian still exists, and the evolution of operators is realized through it. In contrast, our formulation is purely covariant and no longer incorporates a Hamiltonian, since the presence of any Hamiltonian implicitly implies a preferred inertial frame which is a concept that the general relativity aims to discard from the outset. Therefore, the quantum reference system discussed in our article is a general one, where inertial frames hold no special status whatsoever. That is the reason why the general non-inertial effects appear which seem absence in Page-Wootters’ paper. We also observe that it is precisely the quantum fluctuations of the quantum reference system that dictate the non-existence of rigorous inertial frames at the quantum level. Secondly, although both our paper and theirs provide methods for calculating conditional probabilities, these methods are entirely distinct. In their work, the calculation of conditional probabilities is carried out by taking the partial trace via the projection of the quantum state of a specific quantum reference frame subspace which is flat. In other words, this quantum reference frame is still associated with some kind of flat inertial frame, rather than being a general quantum coordinate system. In our framework, conditional probability are computed via projections or partial traces with respect to a generally curved metric of the quantum reference frame subspace, featuring a more natural and covariant geometric construction. Moreover, the entangled state (12) prepared during calibration, is fixed, and does not evolve with respect to external Hamiltonian, instead, the relative evolution of the under-study $|X_{i}\rangle$ with respect to the clock subsystem arises solely from the projection metric $s_{\bar{a}b}\neq\delta_{\bar{a}b}$ of the subspace which only evolves internally. And through geometric method, the paper has found the dynamical equation governing the evolution of the general metric (28), which indirectly gives the time evolution of the conditional probability by the internal evolution of the subspace metric, completely different from Page-Wootters and related theory (PhysRevD.27.2885, ; 1986Quantum, ; Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ). Therefore, this article is not merely a simple re-formulation of the main results of the Page-Wootters framework, rather, it can be regarded as presenting a more general generalization, the new physics is those non-inertial effects coming from the general covariance without a Hamiltonian, just as that general relativity is not a simple re-formulation of Newton’s laws, the new physics is the gravity naturally incorporated.
@@
Although our discussions only focus on the quantum clock (not a quantum spacetime), it is of cause, in essence, a quantum reference frame theory. However, frankly speaking, the results obtained in the existing literature on quantum reference frames (PhysRevD.30.368, ; Rovelli:1990pi, ; RevModPhys.79.555, ; Giacomini:2017zju, ; Loveridge:2017pcv, ; 2020Quant…4..225V, ; Hohn_2020, ; Giacomini:2020ahk, ) are still quite preliminary, in the sense that it is still a long way off from the real goal of this subject: quantum gravity. Compared our works with the existing literature of quantum reference frame, firstly, most current articles on quantum reference frames are still based on the description of quantum reference frames via Hamiltonian. As previously mentioned, this implicitly implies the selection of a preferred inertial frame, making it difficult for these theories to derive non-inertial effects, let alone naturally incorporate gravitational effects. On the other hand, current literature on quantum reference frames still lack a fundamental quantum equivalence principle as a guiding principle to re-interpret the universal quantum properties of material quantum reference frames as the universal properties of the abstract concept of quantum time itself or quantum spacetime itself. In our paper on the quantum equivalence principle (2024arXiv240809630L, ), we first point out that only non-dynamical universal part of quantum second-order moment fluctuation (not coming from its Hamiltonian or mass) of physical reference frames, can universally describe the second-order moments of quantum clocks or quantum spacetime, thereby reconciling the inherent contradiction between the standard Schrödinger equation of quantum mechanics and the equivalence principle (different from most of the literature on the subject, e.g. (Giacomini:2020ahk, ; Das:2023cfu, )). Therefore, our quantum reference frame theory is built upon a covariant theory with a zero Hamiltonian, which is different from most of the literature of the subject. Consequently, in this paper, such quantum clocks naturally give rise to non-inertial effects. Furthermore, we can also discuss here some comparisons between our quantum spacetime reference frame theory (e.g. (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinarxiv, )) beyond the scope of quantum clock and the existing literature on the subject. In our quantum spacetime reference frame (a non-linear sigma model with zero Hamiltonian as a generalization of the quantum clock, which is more suitably treated by functional method rather than the canonical method), the diffeomorphism anomalies naturally lead to the emergence of gravity (Luo:2021zpi, ), and the cancellation of anomalies is connected to the cosmological constant—a feat that has not been truly achieved in the most literature on the quantum reference frame. Secondly, our quantum reference frame theory provides the renormalization group flow of the quantum spacetime (which is, to an approximation, a Ricci flow) and its associated monotonic entropy functional (Luo:2022statistics, ; Luo:2025cer, ), which gives an origin of gravitational entropy, and enabling a global control of a gravitational systems (since a covariant system like gravity lacks a global time coordinate and hence the concept of energy no longer provides a global control like usual dynamical system), a feat not achieved in the current quantum reference frame literature. Moreover, a general quantum coordinate transformation is given by a Jacobian in the functional integral approach which leads to diffeomorphism anomaly and gravity, or equivalently given by an integral transformation via a heat kernel of the conjugate heat flow equations which is coupled to the spacetime Ricci flow.

### v3

@@ V Recovering the Schrödinger Equation
In the previous section, we have seen that the evolution of $|X_{i}\rangle$ or $h_{\bar{i}j}$ w.r.t. the complementary external $t$ (27) is approximately equivalent to the evolution of $\textrm{P}\left(X_{i}|T\right)\overset{separable}{\approx}\textrm{P}\left(X_{i}\right)$ in the separable limit of the entangled state (12), therefore the evolution equations (27) or (28) plays a role analogous to that of the standard Schrödinger equation in describing the evolution of quantum states with respect to absolute external parameters. Therefore, in the following, we will start from the equation (27) and make some approximations to return to the Schrödinger equation limit. First, we linearize the Ricci curvature as $R_{\bar{i}j}\approx-\frac{1}{2}\Delta_{x}h_{\bar{i}j}$, where $\Delta_{x}$ is a linearized the Laplacian-Beltrami operator with respect to $x_{i}$, then eq.(27) becomes
@@
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}=\frac{\partial}{\partial t_{a}}\langle X_{i}|X_{k}\rangle=\frac{\partial\langle X_{i}|}{\partial t_{a}}|X_{k}\rangle+\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}\approx\langle X_{i}|\frac{\partial|X_{k}\rangle}{\partial t_{a}}=e^{-i\omega_{a}t^{a}}\langle X_{i}|\left(\frac{\partial}{\partial t^{a}}-i\omega_{a}\right)|e_{k}\rangle$ (38)
@@
$\frac{\partial h_{\bar{i}k}}{\partial t_{a}}\approx\left(-i\omega^{a}\right)h_{\bar{i}k}$ (39)
@@
To sum up, we have demonstrated that the equation (27) or its linearized approximations (32) or (33) describe the evolution of the metric $h_{\bar{i}j}$ or the basis vectors $|X_{i}\rangle$ in the subspace $\mathscr{H}_{X}$ with respect to the entangled state vector $|T\rangle$ in $\mathscr{H}_{T}$. This is because we observe that the slow-changing part $|e_{i}\rangle$ of the basis vectors $|X_{i}\rangle$ satisfies an approximate “free” Schrödinger evolution equation, where the fast-changing characteristic frequency $\omega^{a}$ plays a role analogous to the mass $M=\hbar\omega^{a}$, in which the mass and the Planck’s constant emerge, and the potential term in the Schrödinger equation is embodied in the connection (scalar extrinsic curvature $K_{a}$) of the covariant derivative.
@@ VI Non-Inertial Effects
where $F_{\bar{a}b}$ reflects a kind of “inertial force” that arises due to the choice of quantum reference system when $|X\rangle$ is considered relative to $|T\rangle$. Under the approximation (35), we return to the situation of an approximately inertial frame where it is inertial-force-free.
@@
in which $G_{\bar{a}b}$ and $\Omega_{\bar{a}b}$ are the real and imaginary part of the metric. As a metric coming from the comparison between $|X^{i}\rangle$ and $|T^{a}\rangle$, it can be interconverted with the previous metrics $h_{\bar{i}j}$ through a tensor coordinate transformation, for example,
$g_{\bar{a}b}=\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X^{j}}{\partial t^{b}}\right\rangle h_{i\bar{j}}=\left\langle\frac{\partial T_{a}}{\partial x_{i}}\right|\left.\frac{\partial T_{b}}{\partial x_{j}}\right\rangle h_{i\bar{j}}$ (48)
@@
$|X_{i}(t+dt)\rangle=|X_{i}(t)\rangle+\left|\frac{\partial X_{i}(t)}{\partial t^{a}}\right\rangle dt^{a}+$... (50)
@@
$2\textrm{Re}\left(d\varTheta\right)$ $=\left|\langle X^{i}(t+dt)|X_{i}(t)\rangle\right|^{2}-1=\textrm{Re}\left\langle\frac{\partial X^{i}}{\partial t^{a}}\right|\left.\frac{\partial X_{i}}{\partial t^{b}}\right\rangle dt^{a*}dt^{b}=G_{\bar{a}b}dt^{a*}dt^{b}$ (51)
@@
Comparing this equation with the definition $\Omega_{\bar{a}b}\equiv g_{\bar{a}b}-g_{\bar{a}b}^{*}$ and by using (49), we see that $\Omega_{\bar{a}b}$ can be interpreted as the Berry curvature (or field strength) $B_{\bar{a}b}$
@@
Next, let us discuss some physical implications of the real and imaginary parts of the non-integrable phase $\varTheta$. We observe that eq.(51) can also be interpreted as an infinitesimal stretching of the length of the state vector $|X\rangle$, which makes the evolution of $|X\rangle$ with respect to $|T\rangle$ no longer unitary in general. Consequently, previously we could only derive the Schrödinger equation with unitary evolution under the assumption of slow-changing of the scalar extrinsic curvature (35) $G_{\bar{a}b}=\textrm{Re}\left(\frac{\partial K_{b}}{\partial t^{a*}}\right)\approx 0$. This is also an inevitable consequence of the relative evolution of subsystems in entangled states, resulting from the quantum superposition (quantum second-moment spreading (Luo2015Dark, ; Luo:2023eqf, ; 2024arXiv240809630L, )) of the $T$ subsystem that serves as the reference frame.
This breakdown of unitarity can also be interpreted as an effect of additional second-order fluctuations or extra broadening. Considering the second-moment spreading of the effective energy $E$ corresponding to the second-moment fluctuations of $t$, namely $\langle\delta E^{2}\rangle=\langle X^{i}|E^{2}|X_{i}\rangle-\langle X^{i}|E|X_{i}\rangle^{2}$, it leads to
$\left|\langle X^{i}(t+dt)|X_{i}(t)\rangle\right|^{2}=1-\langle\delta E^{2}\rangle dt^{2}$ (58)
Therefore, comparing (51) and (58), the real part of the metric $G_{\bar{a}b}$ is related to the second-moment quantum fluctuations of $t$ or the second-moment fluctuations of the effective energy, i.e.
$G_{\bar{a}b}=\langle\delta E_{a}\delta E_{b}\rangle\quad\textrm{or}\quad\left\|G_{\bar{a}b}\right\|\sim\langle\delta E^{2}\rangle$ (59)
where the negative sign of the difference is due to the timelike signature. Hence, this “inertial force” $\left\|\textrm{Re}(F_{\bar{a}b})\right\|\sim\langle\delta E^{2}\rangle$ can also be regarded as arising from the quantum fluctuations of energy or quantum fluctuations of the clock reference frame. We also observe that the breakdown of unitarity caused by the second-order fluctuations of the quantum clock and the broadening of spectral lines (Luo:2023eqf, ) both are coming from the additional “inertial force” effect. Conversely, the effect of the “inertial force” always seems to be associated with changes in the norm of the state vector or the breakdown of quantum unitarity. This non-unitarity resulting from the broadening of the quantum clock or spectral lines will be generalized to more general quantum spacetime reference frames in (Luo:2015pca, ; Luo:2019iby, ; Luo:2021zpi, ; 2024arXiv240809630L, ), and is the fundamental reason of coordinate transformation anomalies, the cosmological constant, and gravity in general quantum spacetime reference frames.
@@
If the phase $\varTheta=\frac{1}{2}\int_{c}dt^{a}K_{a}$ is integrable (independent to the path $c$), meaning that the local Berry curvature $B=B_{\bar{a}b}dt^{a*}\wedge dt^{b}$ is closed ($dB=d^{2}K=0$), and so if this local Berry curvature corresponds to the imaginary part $\Omega_{\bar{a}b}$ of the local metric, the quantum state space equipped with this Hermitian complex metric must be a Kähler space. In this case, the gauge field strength 2-form $\frac{i}{2}B_{\bar{a}b}dt^{a*}\wedge dt^{b}\doteq\frac{i}{2}\Omega_{\bar{a}b}dt^{a*}\wedge dt^{b}$ serves as the Kähler form, reflecting the symplectic structure of the Kähler space. A Kähler space is one that simultaneously possesses a Riemannian structure, a symplectic structure, and is compatible with its complex structure. When the quantum state space is a Kähler space, or equivalently, the phase $\varTheta$ is integrable, due to the enhanced symmetry of the Kähler metric, which can be represented by a single Kähler potential, many problems by solving the multi-component Ricci-flat Kähler-Einstein equations can be reduced to solving a single complex Monge-Ampère equation, which simplifies the problem of solving the Kähler-Einstein equations. When there exists a constant volume form $\sqrt{\det h_{\bar{i}j}}$ of state subspace (w.r.t. another subspace), i.e. a constant Kähler potential $\ln\sqrt{\det h_{\bar{i}j}}$, this degenerates to the case of a unitary quantum mechanics, so the state space of the unitary standard quantum mechanics is Kähler, which is suggested by literature e.g. (Ashtekar:1997ud, ; Bengtsson:2006rfv, ). In particular, when the curvature of its subspace is proportional to the metric, $R_{\bar{a}b}=(N+1)g_{\bar{a}b}$, the quantum state space is homeomorphic to the maximally symmetric sphere $S^{2N+1}$ and this Kähler metric $g_{\bar{a}b}$ is known as the Fubini-Study metric (Pandya:2006mrv, ; Matassa:2020dnc, ). However, the subspace metric $h_{\bar{i}j}$ of interest to us generally does not possess constant curvature or a constant local volume form, so the subspace metric of the quantum state here is generally not Fubini-Study in this framework. Nonetheless, in general, our phase $\varTheta$ is not strictly integrable (depending on the path and history), and the Berry curvature can no longer be locally defined. In this case, the quantum state subspace of the framework is not strictly Kähler but can be regarded as an almost Kähler manifold.
@@ VII Comparison with Some Past Works
Our paper achieves results that are similar in many respects to those in the work by Page and Wootters (PhysRevD.27.2885, ). However, there are also numerous differences between our paper and their study. Firstly, in Page and Wootters’ article, the non-zero Hamiltonian of the clock subsystem still exists, and evolution of the density matrix of the quantum clock is realized through it. In contrast, our formulation is purely covariant and no longer incorporates a Hamiltonian, since the presence of any non-zero Hamiltonian implicitly implies a preferred inertial frame which is a concept that the general relativity aims to discard from the outset. Therefore, the quantum clock reference system discussed in our article is a general one, i.e. any general quantum system can be seen as a quantum clock, where inertial frames hold no special status whatsoever, while Page-Wootters’s theory still presupposes a particular (inertial) type of internal clock. That is the reason why the general non-inertial effects appear which seem absence in Page-Wootters’ theory. We also observe that it is precisely the quantum fluctuations of the quantum reference system that dictate the non-existence of rigorous inertial frames at the quantum level. Secondly, although both our paper and theirs provide methods for calculating conditional probabilities, these methods are entirely distinct. In their work, the calculation of conditional probabilities is carried out by taking the conventional partial trace over the clock (sum over the diagonal elements) which is equivalent to a projection of the quantum state onto a specific quantum clock reference frame subspace which is flat. In other words, this quantum reference frame is still associated with some kind of flat inertial frame, rather than being a general quantum coordinate system. In our framework, conditional probability are computed via partial traces over the clock by contracting by a generally curved metric of the quantum clock reference frame subspace. It is also noted that in Page-Wootters’ calculation of the conditional probability, taking the conventional partial trace and divided by the probability of the quantum clock are two separate steps. However, in our approach, these two steps are synthesized into a single step: the contracting by the clock metric, which features a more natural and covariant geometric construction. Moreover, the entangled state (12) prepared during calibration, is fixed, and does not evolve with respect to external Hamiltonian, instead, the relative evolution of the under-study $|X_{i}\rangle$ with respect to the clock subsystem arises solely from the projection metric $s_{\bar{a}b}\neq\delta_{\bar{a}b}$ of the subspace which only evolves internally. And through geometric method, the paper has found the dynamical equation governing the evolution of the general metric (28), which indirectly gives the time evolution of the conditional probability by the internal evolution of the subspace metric, completely different from Page-Wootters and related theory (PhysRevD.27.2885, ; Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ). And moreover, in this framework, it is more transparent to seen the relation between the conditional probability $\left|\frac{C_{i}}{t_{i}}\right|^{2}$ and the absolution probability $|x_{i}|^{2}$, that is, only in the separable limit of the entangled state (12), or equivalently, the quantum clock can be seen as a global external inertial clock, these two probabilities are equal, which is not easily to seen in the Page-Wootters’ original and subsequence papers. As a consequence, it gives us a deeper understanding to Kuchar’s criticism about the “wrong propagator” in the Page-Wootters’ theory (Kuchar:1991qf, ): that is, the Dirac delta propagator $\delta(x-x^{\prime})\delta(t-t^{\prime})$. The delta propagator is, of course, a correct propagator or conditional amplitude for the general covariant system as a whole without any external evolving parameter (particularly, it is this propagator shows the mass-independence as the quantum equivalence principle (2024arXiv240809630L, ) has pointed out). To crossover to the well-known standard (mass-dependent) propagator, the clock must be separable as an unentangled direct product state, and hence the state of the clock can be factored out as an global external inertial clock in which process a Hamiltonian or mass-dependence emerges in the standard propagator and the conditional probability equals to the standard absolute one. To the best of our knowledge, this resolution to the Kuchar’s criticism is also not explicitly claimed before. For Kuchar’s another criticism of the Page-Wootters mechanism regarding the issue that the conditional probability may violate the Hamiltonian constraint, and more specifically, the problem of whether the clock projection operator be non-commutative with the Hamiltonian constraint operator. The problem also does not exist within our framework. Comparing with the “projection interpretation” of the conditional probability in Page-Wootters’s paper, it is more transparent to see that the “projection” by the clock metric $s_{\bar{a}b}$ is just a “coordinate” transformation (or gauge transformation) of the entangled state (which satisfies the Hamiltonian constraint) in the full Hilbert space, choosing a specific clock metric is just equivalent to choosing a specific “coordinate system”, the conditional probability is a coordinate system fixed and reduced version of invariant inner product. In other words, the clock metric as an induced subspace metric from the full space naturally satisfying the Ricci flat Kähler-Einstein equation of the full space and hence the conditional probability naturally obeys the constraint. It is also for this general covariant reason of the conditional amplitude of the framework, Kuchar’s another criticism of “Inappropriate for the relativistic Klein-Gordon systems” (this criticism is also originated from the non-general-covariance of Page-Wootters’ internal clock time) does not exist in this framework as well, because there must be a “coordinate” transformation to relate the general covariant internal clock time in this framework to an external relativistic Minkovski time for Klein-Gordon systems. To sum up, this article is not merely a simple re-formulation of the Page-Wootters’ framework, rather, it can be regarded as presenting a more general generalization, the new physics are the non-inertial effects coming from the general covariance without a Hamiltonian, just as that general relativity is not a simple re-formulation of Newton’s laws, the new physics is the gravity naturally incorporated.
@@
Although our discussions only focus on the quantum clock (not a quantum spacetime), it is of course, in essence, a quantum reference frame theory. However, frankly speaking, the results obtained in the existing literature on quantum reference frames, (e.g. (PhysRevD.30.368, ; Rovelli:1990pi, ; RevModPhys.79.555, ; Giacomini:2017zju, ; Loveridge:2017pcv, ; 2020Quant…4..225V, ; Hohn_2020, ; Giacomini:2020ahk, ) and references therein) are still quite preliminary, in the sense that it is still a long way off from the real goal of this subject: quantum gravity. Compared our works with the existing literature of quantum reference frame, firstly, most current articles on quantum reference frames are still based on the description of quantum reference frames via non-zero Hamiltonian. As previously mentioned, this implicitly implies the selection of a preferred inertial frame, making it difficult for these theories to derive non-inertial effects, let alone naturally incorporate gravitational effects. And so, the coordinate transformations discussed in these literature can only be seen as transformations between quantum inertial frames, not a general quantum coordinate transformation. On the other hand, current literature on quantum reference frames still lack a fundamental quantum equivalence principle as a guiding principle to re-interpret the universal quantum properties of material quantum reference frames as the universal properties of the abstract concept of quantum time itself or quantum spacetime itself. In our paper on the quantum equivalence principle (2024arXiv240809630L, ), we first point out that only non-dynamical universal part of quantum second-order moment fluctuation (not coming from its Hamiltonian or mass, but coming from a cutoff scale dependent broadening of the Dirac delta propagator mentioned in the previous paragraph) of physical reference frames, can universally describe the second-order moments of quantum clocks or quantum spacetime, thereby reconciling the inherent contradiction between the mass-dependence of the Schrödinger equation of standard quantum mechanics and the equivalence principle (different from most of the literature on the subject, e.g. (Giacomini:2020ahk, ; Das:2023cfu, )). Therefore, our quantum reference frame theory is built upon a covariant theory with a zero Hamiltonian, which is different from most of the literature of the subject. Consequently, in this paper, such internal quantum clocks naturally give rise to non-inertial effects. Furthermore, getting back to our main motivation of the paper: a quantum gravity, (quantum clock discussed in the article is used as a conceptual illustration example), we would like also discuss here some comparisons between our quantum spacetime reference frame theory (e.g. (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinaxiv, )) beyond the scope of quantum clock and the existing literature on the subject. In our quantum spacetime reference frame (a non-linear sigma model with zero Hamiltonian as a generalization of the quantum clock, which is more suitably treated by functional integral approach rather than the canonical relative state approach), the spacetime diffeomorphism anomalies naturally lead to the emergence of gravity (see (Luo:2021zpi, ; 2023AnPhy.45869452L, ; 2024arXiv240809630L, )), and the cancellation of anomalies is connected to the cosmological constant—a feat that has not been truly achieved in the most literature on the quantum reference frame. Secondly, our quantum reference frame theory provides the renormalization group flow of the quantum spacetime (which is, to an approximation, a Ricci flow) and its associated monotonic entropy functional (Luo:2022statistics, ; Luo:2025cer, ), which gives an origin of spacetime and gravitational entropy, and enabling a global control of a gravitational systems (since a covariant system like gravity lacks a global time coordinate and hence the concept of energy/Hamiltonian no longer provides a global control like usual dynamical system), a feat not successfully achieved in most current quantum reference frame literature. Moreover, a general quantum coordinate transformation is given by a Jacobian in the functional integral approach which leads to spacetime diffeomorphism anomaly and gravity, or equivalently, given by an integral transformation via a heat kernel of the conjugate heat flow equations which is coupled to the spacetime Ricci flow. The resulting general quantum coordinate transformation is non-unitary due to the non-inertial effects (including the diffeomorphism anomaly, non-isometric of the Ricci flow and gravity, etc.) coming from the second-order moment quantum fluctuations, which is completely different from most of the existing literature on the subject.